Math 1200 Exam 2 Study Guide

0.0(0)
Studied by 0 people
call kaiCall Kai
learnLearn
examPractice Test
spaced repetitionSpaced Repetition
heart puzzleMatch
flashcardsFlashcards
GameKnowt Play
Card Sorting

1/89

flashcard set

Earn XP

Description and Tags

Contains Chapters 5.1 to 7.2

Last updated 2:11 AM on 10/15/26
Name
Mastery
Learn
Test
Matching
Spaced
Call with Kai
Chat

No analytics yet

Send a link to your students to track their progress

90 Terms

1
New cards

Random

The word _________ suggests an unpredictable result (or outcome).

2
New cards

Simulation

__________ is a technique used to re-create a random event. __________ can be tactile (as in physically flipping a coin) or virtual (using a computer to pretend it is flipping a coin). In both instances, the goal of the __________ is to measure how often a certain outcome (such as a head in coin flipping) is observed.

3
New cards

Random process

A(n) __________ represents scenarios where the outcome of any particular trial of an experiment is unknown, but the proportion (or relative frequency) a particular outcome is observed approaches a specific value as the number of trials increases.


4
New cards

Probability

__________ is a measure of the likelihood of a random phenomenon or chance behavior occurring. __________ deals with experiments that yield random short-term results or outcomes yet reveal long-term predictability.

5
New cards

outcomes

Probability is a measure of the likelihood of a random phenomenon or chance behavior occurring. Probability deals with experiments that yield random short-term results or _________ yet reveal long-term predictability.

6
New cards

Law of Large Numbers

The __________ is often interpreted as a nonexistent law called the Law of Averages by folks who misunderstand the ___________. For example, in baseball you may hear an announcer say that a certain player is due to get a hit because he has gone a number of at-bats without getting a hit.

7
New cards

Experiment

In probability, a(n) _________ is any process with uncertain results that can be repeated. The result of any single trial of the experiment is not known ahead of time. However, the results of the _________ over many trials produce regular patterns that allow accurate predictions.

8
New cards

Sample space

The __________, S, of a probability experiment is the collection of all possible outcomes.

9
New cards

Event

A(n) ________ is any collection of outcomes from a probability experiment. A(n) ________ consists of one outcome or more than one outcome. We will denote ________ with one outcome, sometimes called simple _______, ei. In general, ________are denoted using capital letters such as E.

10
New cards

Fair die

A(n) __________ is one in which each possible outcome is equally likely. For example, rolling a 2 is just as likely as rolling a 5.

11
New cards

Loaded die

We contrast this with a(n) ________, in which a certain outcome is more likely. For example, if rolling a 1 is more likely than rolling a 2,3,4,5, or 6 the die is loaded.

12
New cards

Rules of Probabilities

o     The probability of any event E, P(E),  must be greater than or equal to 0 and less than or equal to 1. That is,  0≤P (E)≤10≤P (E)≤1 .

o     The sum of the probabilities of all outcomes must equal 1. That is, if the sample spaceS={e1,e2,⋯ ,en}S=\left\lbrace{e_1,e_2,\cdots,e_{n}}\rbrace\right.

13
New cards

Rules of Probabilities

What is the picture showing?


14
New cards

probability model

A(n) __________ lists the possible outcomes of a probability experiment and each outcome’s probability. A(n) __________ must satisfy Rules 1 and 2 of the rules of probabilities.

15
New cards

Impossible

If an event is _________, the probability of the event is 0.

16
New cards

Certainty

If an event is a(n) ________, the probability of the event is 1.

17
New cards

Unusual event

A(n) __________ is an event that has a low probability of occurring.

Typically, an event with a probability less than 0.05 (or 5%) is considered _________, but this cutoff point is not set in stone.

18
New cards

Approximation Probabilities Using the Empirical Approach

What is this photo showing?


19
New cards

Equally likely outcomes

An experiment has __________ when each outcome has the same probability of occurring.

20
New cards

Computing Probability Using the Classical Method

What is this picture showing?


21
New cards

Tree diagram

To determine the sample space, construct a(n) __________ to list the equally likely outcomes of the experiment.


22
New cards

Subjective probability

Nonetheless, the forecast of 20% does satisfy the criterion that a probability be between and inclusive. This forecast is known as a(n) ___________. A(n) ____________ of an outcome is a probability obtained on the basis of personal judgment.

23
New cards

Mutually exclusive

Two events are disjoint if they have no outcomes in common. Another name for disjoint events is __________ events.

24
New cards

Venn diagrams

We can use __________ to represent events as circles enclosed in a rectangle. The rectangle represents the sample space, and each circle represents an event.


25
New cards

Addition Rule for Disjoint Events

These pictures are showing?


26
New cards

The General Addition Rule

What is this picture showing?

<p>What is this picture showing?</p>
27
New cards

Contingency table or two-way table

This table is called a(n) ___________, because it relates two categories of data. The row variable is marital status, because each row in the table describes the marital status of an individual. The column variable is gender. Each box inside the table is called a cell.


28
New cards

Complement of an Event

Let S denote the sample space of a probability experiment and let E denote an event. The complement of E denoted EC, is all outcomes in the sample space S that are not outcomes in the event E.


29
New cards

Complement Rule

What are these pictures showing?


30
New cards

Independent

Two events E and F are __________ if the occurrence of event E in a probability experiment does not affect the probability of event F.

31
New cards

Dependent

Two events are _________ if the occurrence of event E in a probability experiment affects the probability of event F.

32
New cards

True

(True or False) Two events that are disjoint are not independent.

33
New cards

Multiplication Rule for Independent Events

What is this picture showing?


34
New cards

Multiplication Rule for n Independent Events

What is this picture showing?


35
New cards

Conditional Probability

The notation P(F|E) is read “the probability of event given event F given event E.” It is the probability that event F occurs, given that event E has occurred.

36
New cards

Conditional Probability Rule

What is this picture showing?


37
New cards

General Multiplication Rule

What is this picture showing?


38
New cards

Multiplication Rule of Counting

What is this picture showing?


39
New cards

Factorial symbol

If n ≥ 0 is an integer, the ________, 𝑛!, is defined as follows:


n! = n(n−1) ⋅ ⋯ ⋅ 3 ⋅ 2 ⋅ 1

0!=1     1!=1

40
New cards

Permutation

A(n) __________ is an ordered arrangement in which r objects are chosen from n distinct (different) objects so thatr≤nr≤n and repetition is not allowed. The symbol nPr_{n}P_{r}, represents the number of _________s of r objects selected from n objects.


41
New cards

Number of Permutations of n Distinct Objects Taken r at a Time

What is this picture showing?


42
New cards

Combination

A(n) ___________ is a collection, without regard to order, in which r objects are chosen from n distinct objects with r ≤ n and without repetition. The symbol nCr represents the number of combinations of n distinct objects taken r at a time.


43
New cards

Number of Combinations of n Distinct Objects Taken r at a Time

What is this picture showing?


44
New cards

Permutations with Nondistinct Items

What is this picture showing?


45
New cards

Use classical approach: P(E) = N(E) / S(S)

Which method do I use?

Are you finding the probability of a single event? - Yes

Are the outcomes equally likely? - Yes

46
New cards

Use P(E) = Frequency of E / Number of trials

Which method do I use?

Are you finding the probability of a single event? - Yes

Are the outcomes equally likely? - No

Is it possible to get empirical data? - Yes

47
New cards

Use subjective probability

Which method do I use?

Are you finding the probability of a single event? - Yes

Are the outcomes equally likely? - No

Is it possible to get empirical data? - No

48
New cards

Use P(E or F) = P(E) + P(F)

Which method do I use?

Are you finding the probability of a single event? - No

Do the events involve ‘OR‘, ‘AND‘, or ‘At Least‘? - OR

Are the events disjoint (mutually exclusive)? - Yes

49
New cards

Use P(E or F) = P(E) + P(F) - P(E and F)

Which method do I use?

Are you finding the probability of a single event? - No

Do the events involve ‘OR‘, ‘AND‘, or ‘At Least‘? - OR

Are the events disjoint (mutually exclusive)? - No

50
New cards

Use P(E and F) = P(E) * P(F)

Which method do I use?

Are you finding the probability of a single event? - No

Do the events involve ‘OR‘, ‘AND‘, or ‘At Least‘? - AND

Are the events independent? - Yes

51
New cards

Use P(E and F) = P(E) * P(F | E)

Which method do I use?

Are you finding the probability of a single event? - No

Do the events involve ‘OR‘, ‘AND‘, or ‘At Least‘? - AND

Are the events independent? - No

52
New cards

Use P(at least 1) = 1 - P(none)

Which method do I use?

Are you finding the probability of a single event? - No

Do the events involve ‘OR‘, ‘AND‘, or ‘At Least‘? - At Least

53
New cards

Use the Multiplication Rule of Counting (p * q * r * …)

Which method do I use?

Are you making a sequence of choices? - Yes

Are the number of choices at each stage independent of previous choices - Yes

54
New cards

Use tree diagram

Which method do I use?

Are you making a sequence of choices? - Yes

Are the number of choices at each stage independent of previous choices - No

55
New cards

Use nPr = n! / (n - r)!

Which method do I use?

Are you making a sequence of choices? - No

Does the order of arrangements matter? - Yes

Are you selecting all n objects, or only r objects? - Only r

56
New cards

Use n!

Which method do I use?

Are you making a sequence of choices? - No

Does the order of arrangements matter? - Yes

Are you selecting all n objects, or only r objects? - Only n

Are the objects all distinct? - Yes

57
New cards

Use n! / n1! n2! … nk!

Which method do I use?

Are you making a sequence of choices? - No

Does the order of arrangements matter? - Yes

Are you selecting all n objects, or only r objects? - Only n

Are the objects all distinct? - No

58
New cards

Use nCr = n! / r!(n - r)!

Which method do I use?

Are you making a sequence of choices? - No

Does the order of arrangements matter? - No

59
New cards

Random variable

A(n) __________ is a numerical measure of the outcome of a probability experiment, so its value is determined by chance. __________ are typically denoted using capital letters such as X.

60
New cards

Discrete random variable

A(n) __________ has either a finite or countable number of values. The values of a __________ can be plotted on a number line with space between each point.

61
New cards

Continuous random variable

A(n) __________ has infinitely many values. The values of a __________ can be plotted on a line in an uninterrupted fashion.

62
New cards

Probability distribution

The __________ of a discrete random variable provides the possible values of the random variable and their corresponding probabilities. A(n) _________ can be in the form of a table, graph, or mathematical formula.

63
New cards

Rules for a Discrete Probability Distribution

What is this picture showing?


64
New cards

The Mean of a Discrete Random Variable

What is this equation?


<p>What is this equation?</p><p></p>
65
New cards

Expected value

Because the mean of a random variable represents what we would expect to happen in the long run, it is also called the __________, E(X). The interpretation of __________ is the same as the interpretation of the mean of a discrete random variable.

66
New cards

Standard Deviation of a Discrete Random Variable

What is this equation?

<p>What is this equation?</p>
67
New cards

Binomial probability distribution

The __________ is a discrete probability distribution that describes probabilities for experiments in which there are two mutually exclusive (disjoint) outcomes. These two outcomes are generally referred to as success (such as making a free throw) and failure (such as missing a free throw).

68
New cards

Binomial experiment

An experiment is said to be a(n) __________ if

o   The experiment is performed a fixed number of times. Each repetition of the experiment is called a trial.

o   The trials are independent. This means that the outcome of one trial will not affect the outcome of the other trials.

o   For each trial, there are two mutually exclusive (disjoint) outcomes: success or failure.

o   The probability of success is the same for each trial of the experiment.

69
New cards

Binomial random variable

Let the random variable be the number of successes in trials of a binomial experiment. Then is called a(n) _________.

70
New cards

Binomial Probability Distribution Function

What is this picture showing?


71
New cards

≥\ge

What math symbol uses the phase

“at least“ or “no less than“ or “greater than or equal to“

72
New cards

>>

What math symbol uses the phase

“more than“ or “greater than“

73
New cards

<<

What math symbol uses the phase

“fewer than“ or “less than“

74
New cards

≤\le

What math symbol uses the phase

“no more than“ or “at most“ or “less than or equal to“

75
New cards

=

What math symbol uses the phase

“exactly“ or “equals“ or “is“

76
New cards

Mean (or Expected Value) and Standard Deviation of a Binomial Random Variable

What is this picture showing?


77
New cards

Uniform probability distribution

The random variable X can be any value in the interval from to that is, 0≤x≤300\le x\le30 . Because any two intervals of equal length between 0 and 30 inclusive, are equally likely, the random variable X is said to follow a(n) __________.

78
New cards

Probability density function (pdf)

A(n) __________ is an equation used to compute probabilities of continuous random variables. It must satisfy the following two properties:

  • The total area under the graph of the equation over all possible values of the random variable must equal

  • The height of the graph of the equation must be greater than or equal to for all possible values of the random variable.


79
New cards

Model

In mathematics, a(n) _________ is an equation, table, or graph used to describe reality.

80
New cards

Normal curve

The red curve in a model called the _________, which is used to describe continuous random variables that are said to be normally distributed.


81
New cards

Normal probability distribution

A continuous random variable is normally distributed, or has a ___________, if its relative frequency histogram has the shape of a normal curve.

82
New cards

Inflection points

The points at x=μ−σx=\mu-\sigma and x=μ+σx=\mu+\sigma are the __________ on the normal curve, the points on the curve where the curvature of the graph changes.

83
New cards

Properties of the Normal Density Curve

o     The normal curve is symmetric about its mean, μ\mu .

o    Because mean = median = mode, the normal curve has a single peak and the highest point occurs at x=μx=\mu .

o     The normal curve has inflection points at μ−σ\mu-\sigma and μ+σ\mu+\sigma .

o     The area under the normal curve is 1

o     The area under the normal curve to the right of μ\mu equals the area under the curve to the left of μ\mu , which equals 12\frac12 .

o     As x increases without bound (gets larger and larger), the graph approaches, but never reaches, the horizontal axis. As x decreases without bound (gets more and more negative), the graph approaches, but never reaches, the horizontal axis.

o   The Empirical Rule:

  • Approximately 68% of the area under the normal curve is between x=μ−σx=\mu-\sigma and x=μ+σx=\mu+\sigma ;

  • approximately 95% of the area is between x=μ−2σx=\mu-2\sigma and x=μ+2σx=\mu+2\sigma ;

  • approximately 99.7$ of the area is between x=μ−3σx=\mu-3\sigma and x=μ+3σx=\mu+3\sigma .


84
New cards

The Empirical Rule

What is this picture showing?


85
New cards

Normal probability density function

The equation (or model) used to determine the probability of a continuous random variable is called a probability density function (or pdf). The ___________ is given by


86
New cards

Area Under a Normal Curve

Suppose that a random variable is normally distributed with mean and standard deviation . The area under the normal curve for any interval of values of the random variable represents either

o   the proportion of the population with the characteristic described by the interval of values or

o   the probability that a randomly selected individual from the population will have the characteristic described by the interval of values.


87
New cards

Standardizing a Normal Random Variable

What is this picture showing?


88
New cards

Standard normal random variable

If a normal random variable X has mean different from 0 or a standard deviation different from 1, X can be transformed into a(n) ___________ Z whose mean is 0 and standard deviation is 1.

89
New cards

Standard normal curve

What do you call the graph shown in Figure 10?


90
New cards

za

The notation __ (pronounced “z sub alpha”) is the z-score such that the area under the standard normal curve to the right of __ is a.